8 May 2015

Also today science and packing down

 Nanna and Anna are passing while doing radar and GPS work.

Nanna and Anna are passing while doing radar and GPS work.

As yesterday, people in camp were doing two lines of work: Collecting samples and doing radar measurements (science) and packing down in camp (logistics).

As the wind dropped from 7 m/s to 3 m/s this afternoon, it was time to remove the cover on the carpenters garage. Also, all markers on the ski landing area were retrieved. They will be setup again at EGRIP.

After Pat had arranged fuel bladders on the GrIT carpet sled, some 19,000 liter fuel was transferred from NEEM tanks to the bladders. It is much easier to transport fuel in bladders than in tanks.

Dorthe is troubleshooting the control for the winch that will be used to lower our borehole logging device all the way to the bottom of the 2538 m NEEM borehole.

We have had contact with our colleagues on Renland ice cap, and we were happy to learn that they now are fully deployed on the ice cap and that camp construction is going fine.

What we have done today:

1.    Maintenance and adjustment of the GrIT sled.
2.    Transferring fuel from NEEM tanks 2 and 3 to GrIT bladders.
3.    Setting up equipment for borehole logging. Troubleshooting winch control.
4.    Continuing pit study and hand augering of shallow cores.
5.    Practice run operating the GPR radar and GPS.
6.    Main tarp removed from carpenters garage.
7.    Removed all ski landing area markers.
8.    Surface snow sampling.

Ad.2: More than 5000 liter remain in NEEM tank 1

Weather: Mostly thin, broken cloud cover. Temp. – 19°C to -26°C, 3-7 m/s from S and SW. Visibility: Mostly 1-3 km. Few light snow showers and slight fog.

FL, J.P. Steffensen

Paul at the pit study site with halo in the background.

← Previous entry   Next entry →